OpenStudy (anonymous):

if f(x) = 3x-2 and g(x)= x^2 + 1 find (f + g) (x) answers x^2 + 3x + 1 x^2 + 3x-1 4x^2 - 1 2x + 3

OpenStudy (cwrw238):

just add the 2 functions 3x - 2 + x^2 + 1 =

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

what do you mean by add the second functions? Like add the 3x-2 + x^2 + 1? or what? @cwrw238

OpenStudy (turingtest):

not add the second functions, add the two functions i think you added them wrong, look at the constant terms

OpenStudy (cwrw238):

you can only add like terms ;- x^2 + 3x cannot be added but you can add the 2 constants

OpenStudy (anonymous):

so the 2 and the 1 ?

OpenStudy (cwrw238):

the -2 and the 1 - don't forget the sign!

OpenStudy (anonymous):

B? because the answer comes to -1..

OpenStudy (triciaal):

(f + g) (x) = f(x) + g(x) = (3x -2) + (x^2 + 1) like @cwrw238 said you can only add like terms

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

ok thanks!!

OpenStudy (triciaal):

you are welcome